# SparQ Attention: Bandwidth-Efficient LLM Inference

### v4

@@ 6 Benchmarking
The results above use a theoretical cost model of total memory transfers (the number of scalar elements transferred to and from memory per token), allowing us to evaluate SparQ Attention independently of the specific hardware setup and number formats used. To validate our findings, we performed a set of microbenchmarks of an attention operation in isolation.
@@
Table 4: GPU performance with batch size $64$, sequence length $S=4096$, $r=32$, $k=128$. The theoretical speed-up is 6.4×6.4$\times$6.4 ×.
@@ Acknowledgements
We would like to thank Daniel Justus, Paul Balança and Andrew Fitzgibbon for their helpful input and feedback on this work.
@@ Appendix A Detailed Results
Figure A4: Heatmaps of the results for the needle in a haystack task, as outlined in Dhinakaran (2024). For various sequence lengths (comprising of essays by Paul Graham), the needle (which is the sequence ““The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day”) is inserted into the sequence at various depths, and the model is prompted to answer what the best thing to do in San Francisco is. The heatmaps show whether the model returns the correct (blue) or incorrect (red) answer. We evaluated this task over SparQ Attention, H2O and LM-Infinite (in addition to the dense baseline) on the togethercomputer/LLaMA-2-7B-32K model, for two compression ratios.

### v5

@@ 6 Benchmarking
The results above use a theoretical cost model of total memory transfers (the number of scalar elements transferred to and from memory per token), allowing us to evaluate SparQ Attention independently of the specific hardware setup and number formats used. To validate our findings, we performed a set of microbenchmarks of an attention operation in isolation, in addition to end-to-end performance benchmarks.
@@ 6.1 Microbenchmarks
Table 4: GPU microbenchmark performance with batch size $64$, sequence length $S=4096$, $r=32$, $k=128$. The theoretical speed-up is 6.4×6.4$\times$6.4 ×.
@@
6.2 End-to-End Performance
In addition to the positive microbenchmark results, we further highlight the practical improvements SparQ Attention offers by benchmarking performance of the entire Transformer model. We implemented SparQ Attention in llama.cpp, and evaluated its performance on AMD EPYC CPU systems with up to $256$GB memory. We measured the time it took to generate a single token, given an existing sequence length $S$.
The results, as seen in Figure 9, show that SparQ Attention attains speedups at all sequence lengths evaluated, compared to the dense baseline. At the longest sequence lengths considered, SparQ Attention achieves a 2.5×2.5$\times$2.5 × speedup, showcasing the benefits of reducing the data transfer associated with attention.
Figure 9: End-to-end speedup results for SparQ Attention compared to the dense baseline with Llama $2$ $7$B. This was achieved with batch size $1$ and a compression ratio of $1/8$ on CPU, with model weights represented in $8$-bits and the KV cache represented in $16$-bits. The $\mathcal{M}_{\mathrm{dense}}/\mathcal{M}_{\mathrm{SparQ}}$ line is a theoretical result, illustrating the upper bound of speedups attained with SparQ Attention.
@@ Acknowledgements
We would like to thank Daniel Justus, Paul Balança, Oscar Key and Andrew Fitzgibbon for their helpful input and feedback on this work.
@@ Appendix A Detailed Results
Figure A4: Heatmaps of the results for the needle in a haystack task, as outlined in Dhinakaran (2024). For various sequence lengths (comprising of essays by Paul Graham), the needle (which is the sequence “The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day”) is inserted into the sequence at various depths, and the model is prompted to answer what the best thing to do in San Francisco is. The heatmaps show whether the model returns the correct (blue) or incorrect (red) answer. We evaluated this task over SparQ Attention, H2O and LM-Infinite (in addition to the dense baseline) on the togethercomputer/LLaMA-2-7B-32K model, for two compression ratios.